Home‎ > ‎Reports‎ > ‎Crew 108‎ > ‎

day03


Crew #108 Log Book for December 05, 2011

Photos of the day





Commander's Check-In Report

Date and time: 

5 December, 4 pm


Written by: 

Charlotte Poupon, Crew 108 Commander


Crew Physical Status: 

Quite Good.


Time Departed/Returned from EVA:  

10:00am to 11.45 am


Brief Narrative of Field Mission Results: 

N/A


EVA Data/Interpretations: 

Difficult to communicate with the radio.

Two EVAs of 3 people each was organized, 3 others assuming the HabCom.
EVAs will be documented in later report.


Engineering/Hab Maintenance:

For an unknown reason, on the morning, the Commander's room was locked
after her. We had to unscrew the locker, and replace it after opening
the opening. The locker is now 100% operational without any more
trouble.


Report Transmission Schedule:

Engineering + Journalist + EVA reports, due at 8 pm


Plans for Tomorrow: Receive a film-maker 


Inventory:

Unfortunately, yesterday night, at diner, the glassed brownie mold was
broken after a temperature shock between stove and sink. No one was
wounded. We trash the glass parts in a box.


Support Requested:

Is there a way to get into the Musk Observatory?


Miscellaneous: 

It was our first cooking day.

For breakfast, we could enjoy bread with jam, Nutella or peanut butter.

At lunch we had a soup with one pancake and everybody seams to be satisfied.

For diner, our Chef will be creative.

We'll make bread for tomorrow as it is still a cooking day.

Science Report

Date:

5 December 2011


Written by: 

Usha Lingappa, Crew 108, Astrobiologist 


Text:

On the third EVA, we found a small pond that was entirely frozen over. Using equipment sterilized with ethanol in the field, shavings from the ice and mud from beneath the ice were aseptically sampled into sterile whirl-pak sample bags.

Upon return to the Hab, the mud and water samples were plated onto Luria Bertani agar plates, and placed in the incubator (fluctuating between 20 and 30ºC). I intend to isolate organisms from the pond water and mud samples, which can serve as analogs for Martian ponds (or polar ice caps), and underlying mud.

Engineering Report

Date and Time:

5 December 2011, 17:00


Generators/ Electricity
Kitty
Hobbs
Coolant
Oil
Battery Voltage
Casper Battery Voltage

Honey

Fuel
Diesel Tank 1/2
Propane Tank 10%
Gasoline (gallons) 7
Oil Available (quarts) Unknown
Notes:

Inverter/ Batteries
Heat Tapes Operating 24/7
One Meter
Amps Where is this?
Volts Where is this?
Watts Where is this?
kWHr Where is this?
Vehicles
New Blue
Used no
Oil not checked
Tyres good
Odometer (miles) not checked

ATV 1
Gasoline used (gallons) 1
Oil Checked yes
Tyres good

ATV 2
Gasoline used (gallons) 1
Oil Checked yes
Tyres good

ATV 3
Gasoline used (gallons) 1
Oil Checked yes
Tyres good

Water
Trailer (inches from bottom) 3
Outside Potable (inches from top) 8
Inside Potable (markings) Top Green Line

Temperature (deg F)
Min Don't know
Current Is there an external temperature sensor?
Max Don't know
Humidity (%)
Min Don't know
Current 10
Max Don't know

Observatory
radio telescope not used
Musk observatory not used
Hab
HVAC OK
Mars Surface Suits OK except for loose joints on 4 and 5
Computers/Network Ok except webcams not uploading
Appliances OK
GreenHab Offline
Observatory not used

Journalist Report

Date: 

5 December 2011


Writen by: 

Alicia Framis, Crew 108, Journalist


Text:

Today Marti, the mouse was transported in the middle of the dessert by 3 astronauts. I was awaiting the exiting news. Usha too, wanted to
take some examples of bacteria …here in the Hab two astronauts were trying to find the way to reach more water. I prepared a delicious tea.

In a few hours, we received some more news from the EVA team, Marti does not want to leave the box, however it is its destiny to be free. We had a nice soup together, sharing food is a good ritual to stay together. 

After all that action we went outside for some more action. In the middle of the dessert with only our suits on, felling the breath inside our helmets, around us a nothing but silence, just like on Mars.

EVA 3 Report

Date and time:

5 December 2011


Written by: 

Charlotte Poupon, Crew 108, Commander


Time:

10.00 to 11.45 am


Crew Members:

Charlotte, Ashley and Usha


Site Location:

North of Hab


Transit Mode:

ATV


Objectives:

This EVA was a mousitarian one since we released the mouse trapped the night before in the kitchen Hab.


Results:

We went mainly north and found a nice place to release him (her) with nice desert vegetation.

The mouse was pretty shy to get out of the plastic box we put in but after a couple of minutes, we managed to free him. He jumped from plant to plant and we went back to our ATV.

Just when we were about to go, the mouse freakishly charged us. We put the
plastic box on the floor and the mouse decided to step in it !

The EVA was an ethic dilemma… to leave or not to leave a mouse that loves human company? So we left the box open on the floor and went on our EVA… deciding that once back, if the mouse is still there will take it back to the Hab… but the mouse had disappeared and we took back the empty plastic box to the Hab.

During this EVA, the astrobiologist also took some samples of mud and ice from a tiny lake, and rocks. She will write a scientific report byt 10 pm.

Lessons Learned:

EVA 4 Report

Date and time:

5 December 2011


Written by:


Michael LeClair, Crew 108, XO



Time:

10:15am to 11:30 am


Crew Members:

Lotto (Engineer, EVA lead), LeClair (Repairman)


Site Location:

Area surrounding the Hab


Transit Mode:

Foot


Objectives:

Continue work on pressurized tunnel


Results:

Almost completed pressurized tunnel


Lessons Learned:

None

EVA 5 Report

Date and time:

5 December 2011


Written by:


Ashley S. Dale, Crew 108, Engineer



Time:

1:20pm to 3:35 pm


Crew Members:

LeClair (Sortie Lead, Geologist), Lotto (Engineer, Geo-tech), Framis (Journalist)


Site Location:

See Results


Transit Mode:

ATV


Objectives:

This EVA was a mousitarian one since we released the mouse trapped the night before in the kitchen Hab.


Results:

Start point 100m NW of Intersection of Lowell Highway and Cactus Road. Crew joined road at western valley cut opening, continuing through to opposite opening.

First Stop: At mid-point, Lotto initiated a valley wall climb onto very large collapsed boulder Morrison unit, resulting in a now documented find. Three possible concretion specimens were gathered. Best candidate a 4 by 4 by 5 cm spheroid, brown outer layer, light grey inner layer coloured sample with one pronounced internal zoning.

Second Stop: EVA party climbed Mount Sagewood.

Sortie party continued on to scout the mouth Candor Chasma via Chapman Way with secondary objective of testing vehicle range and fuel burn.


Transportation Units and Environmental Suits nominal

Possible Cyanobacteria found


Lessons Learned:

External Radio Communications highly unreliable